# Fast Inference Using Automatic Differentiation and Neural Transport in Astroparticle Physics

### v1

@@ Abstract
Multi-dimensional parameter spaces are commonly encountered in astroparticle physics theories that attempt to capture novel phenomena. However, they often possess complicated posterior geometries that are expensive to traverse using techniques traditional to this community. Effectively sampling these spaces is crucial to bridge the gap between experiment and theory. Several recent innovations, which are only beginning to make their way into this field, have made navigating such complex posteriors possible. These include GPU acceleration, automatic differentiation, and neural-network-guided reparameterization. We apply these advancements to astroparticle physics experimental results in the context of novel neutrino physics and benchmark their performances against traditional nested sampling techniques. Compared to nested sampling alone, we find that these techniques increase performance for both nested sampling and Hamiltonian Monte Carlo, accelerating inference by factors of $\mathord{\sim}100$ and $\mathord{\sim}60$, respectively. As nested sampling also evaluates the Bayesian evidence, these advancements can be exploited to improve model comparison performance while retaining compatibility with existing implementations that are widely used in the natural sciences.
**footnotetext: Equal contribution.
@@ 1 Introduction
The Standard Model of particle physics, our best description of the fundamental building blocks of the Universe, is an incomplete description of nature. To better capture reality, theories that go beyond the Standard Model are needed Murayama:2007ek ; Allanach:2016yth ; Lee:2019zbu . These theories often feature multi-dimensional parameter spaces (with order $10$ parameters or more) that must be explored to determine regions that are allowed by the observed data Fan:2010gt ; Feroz:2011bj ; Levi:2018nxp ; Arguelles:2022lzs ; Isidori:2023pyp ; Amaral:2023tbs . However, methods for doing so that are traditional to the physics community, such as nested sampling (NS), often slow down when navigating complicated posterior geometries 2018arXiv180306387C ; Cerdeno:2018bty ; 2020arXiv201215286A ; Trotta2008TheIO . This problem is particularly acute in the limit-setting regime, where no significant novel signal is observed and the data is used to restrict the domain of the theory. In this regime, the likelihood is often diffuse and lacks strong modes, rendering it difficult to explore. The problem is exacerbated when the likelihood is expensive to evaluate, as is often true due to the complexity of theoretical frameworks. These issues make it challenging to diagnose convergence failures, incorporate new experimental data, and produce timely studies.
These problems can be tackled using a combination of methods. Leveraging recent powerful numerical computing frameworks, such as JAX jax2018github , and probabilistic programming frameworks, such as numpyro bingham2018pyro ; phan2019composable , allows us to automatically differentiate likelihood functions and thus sample efficiently using Hamiltonian Monte Carlo (HMC) Duane:1987de ; Neal2011MCMCUH and the No-U-Turn Sampler (NUTS) hoffman2014no . For likelihood functions that are computationally expensive but vectorizable, GPU acceleration can also confer a significant performance boost. Finally, normalizing flows can be used as a reversible transformation to obtain a posterior distribution that is easier to explore in a technique known as neural transport (NeuTra) 2019arXiv190303704H .
@@
2 Background
2.1 Sampling Methods
Nested sampling
Nested sampling is a family of Monte Carlo methods for calculating integrals over a parameter space skilling2004 ; Buchner2023 . It can also be used for Bayesian computation and has gained popularity for its success in multi-modal problems and its self-tuning feature skilling2006 . NS has many applications in various fields, such as biology russel2019model ; arvindekar2024optimizing , linguistics sagart2019dated ; robbeets2018bayesian and the physical sciences ashton2022nested —including particle physics Yallup:2022yxe ; Dutta:2020che .
@@ 2.1 Sampling Methods
$\mathcal{Z}\equiv\int p(\boldsymbol{\theta}|\boldsymbol{X})\mathop{}\!\mathrm{d}\boldsymbol{\theta}\simeq\sum_{i=1}^{N_{\text{iter}}}\Delta V_{i}L_{i}+\frac{V_{\text{end}}}{N}\sum_{j=1}^{N}L_{j},$ (2)
@@
HMC/NUTS
Hamiltonian Monte Carlo is a powerful sampling algorithm that leverages gradient information to efficiently explore complex probability distributions Duane:1987de ; Neal2011MCMCUH ; Betancourt:2017ebh . HMC offers improved performance compared to traditional Markov chain Monte Carlo (MCMC) methods using a simulation of Hamiltonian dynamics to explore a target parameter space more efficiently than random walks. However, The effectiveness of HMC relies on careful tuning. The NUTS sampler with dual-averaging alleviates this need by adaptively adjusting HMC parameters pmlr-v130-hoffman21a ; pmlr-v151-hoffman22a , and it has been widely adopted and implemented across various probabilistic programming frameworks hoffman2014no ; bingham2018pyro ; phan2019composable ; carpenter2017stan ; pymc2023 .
HMC is extensively used across both the natural and social sciences, and it has been recently employed for analyzing neutrino experiment results Duane:1987de ; Freedman:2021ahq ; sharma2021understanding ; Wang2022BrainPyAF ; Fazio2024GaussianDS ; NOvA:2023iam . Because HMC is a gradient-based sampling method, it requires the model to either support automatic differentiation or have known derivatives. Interest in it has thus led to new implementations of models using frameworks compatible with automatic differentiation Campagne:2023ter ; Ruiz-Zapatero:2023hdf .
Neural Transport
While NS and HMC are better at exploring complex posteriors than simpler methods, such as basic rejection sampling with uniform proposals, sampling efficiency still drops for sufficiently challenging posteriors in multiple dimensions Buchner2023 ; Papaspiliopoulos2007AGF . Under these circumstances, reparameterization can improve the performance of inference algorithms pmlr-v119-gorinova20a , which can be non-trivial for likelihoods that cannot be easily expressed as hierarchical Bayesian models.
One way to remedy this is to use a reversible transformation to map a posterior distribution onto one that is easier to sample, such as a standard normal distribution 2019arXiv190303704H . Instead of directly sampling from the posterior $p(\boldsymbol{\theta}|\boldsymbol{X})$ using HMC or NS, we can consider a bijective function $\boldsymbol{f}$ that defines the reparameterization of our distribution. This function acts on the reparameterized latent space $\boldsymbol{z}$, such that $\boldsymbol{\theta}=\boldsymbol{f}(\boldsymbol{z})$. We can then instead sample from the distribution
@@
where $|\partial\boldsymbol{f}/\partial\boldsymbol{z}|$ is the determinant of the Jacobian of the parameter transformation. Eq. 3 generally follows from the reparameterization of probability distributions, as shown in Appendix A.
Subsequently, the pushforward $\boldsymbol{f}(\boldsymbol{z})$ can be used to obtain the desired samples of $p(\boldsymbol{\theta}|\boldsymbol{X})$. Normalizing flows are a good candidate to parameterize the bijective function, as they can be highly expressive and can be fit by minimizing the Kullback–Leibler (KL) divergence. Using them to reparameterize Bayesian models is known as Neural Transport (NeuTra), as introduced in 2019arXiv190303704H . There, NeuTra with inverse autoregressive flows was shown to improve sampling performance in problems with difficult posterior distributions. Instead of inverse autoregressive flows, we use block neural autoregressive flows, which demonstrate improved performance over earlier methods 2019arXiv190404676D .
To train a normalizing flow that can transform the posterior distribution of interest into a standard normal distribution, we maximize the evidence lower bound (ELBO) 2019arXiv190303704H ; Ranganath2013BlackBV , which is equivalent to minimizing the KL divergence kingma2013auto . Given $\boldsymbol{f}_{\phi}(\boldsymbol{z})$—the transformation represented by a normalizing flow with parameters $\phi$—the distribution over $\boldsymbol{\theta}$ is
@@
2.2 Neutrinos, Astroparticle Experiments, and Novel Physics
Neutrinos at Astroparticle Experiments
Astroparticle Detector Solar Neutrino ($\nu$) Target Atom ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\nu}{\color[rgb]{1,1,1}{}_{\alpha}}$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\nu}{\color[rgb]{1,1,1}{}_{\beta}}$ (Not to scale) Sun
@@ 2.2 Neutrinos, Astroparticle Experiments, and Novel Physics
[…]
Neutrinos are fundamental particles created by the radioactive decay of unstable atomic nuclei. This occurs in places like nuclear reactors, atomic bombs, and the Sun. Neutrinos come in three flavors—electron ($e$), muon ($\mu$), and tau ($\tau$)—and quantum mechanical effects lead to them oscillating between these flavors Davis:1968cp ; SNO:2001kpb . This transmutation is one of the best pieces of evidence we have for physics beyond the Standard Model, making them excellent candidates with which to search for novel phenomena  SajjadAthar:2021prg . Experiments sensitive to neutrino interactions can thus probe phenomena beyond the Standard Model.
Among these experiments are direct detection experiments. Originally leading the hunt for dark matter, these astroparticle detectors are beginning to have exquisite sensitivity to neutrinos produced in the Sun Aalbers:2022dzr . Two such experiments are XENON1T XENON:2017lvq and PandaX-4T PandaX:2018wtu . Solar neutrinos travel to these detectors, interact with the material, and cause the recoil of either a target nucleus or electron. We illustrate this in Fig. 1 (left). We can count the number of nuclear recoils (NRs) and/or electron recoils (ERs) and use it to answer the question of whether or not it is consistent with the prediction from the Standard Model of particle physics.
Neutrino Non-Standard Interactions
@@
Neutrino physics that goes beyond the Standard Model can be described using the framework of neutrino non-standard interactions Wolfenstein:1977ue ; Guzzo:1991hi ; Guzzo:1991cp ; Gonzalez-Garcia:1998ryc ; Guzzo:2000kx ; Guzzo:2001mi ; Gonzalez-Garcia:2004pka . In this framework, we can make predictions by sampling from an $8$-dimensional parameter space defined by the coordinates $(\varepsilon_{\alpha\beta},\eta,\varphi)$ Amaral:2023tbs . Here, the $6$ parameters $\varepsilon_{\alpha\beta}$ characterize the overall strength and type of interaction. There is one such parameter for each flavor combination $\alpha\beta$ ($ee$, $e\mu$, $e\tau$, etc.), with equal flavor indices meaning that the neutrino type remains the same after the interaction and unequal indices indicating that it changes. The remaining two parameters $(\eta,\varphi)$ are angles describing how strongly the new interaction takes place with protons, neutrons, and electrons. The angle $\eta$ models how strongly the interaction is with the neutron, and the angle $\varphi$ characterizes the strength of the interaction between the electrically charged proton and electron. We visualize the NSI parametrization in Fig. 1 (right) and further discuss it in Appendix B.
@@
Within this framework, we calculate the number of expected NRs and ERs using the SNuDD Python package SNuDD-code ; Amaral:2023tbs . Given a choice of NSI parameters, it computes the number of recoils
$N_{\mathrm{NR/ER}}=\iint\texttt{SNuDD}(\varepsilon_{\alpha\beta},\eta,\varphi;E_{\nu},E_{R})\mathop{}\!\mathrm{d}E_{\nu}\mathop{}\!\mathrm{d}E_{R}\,,$ (6)
where the integrand is a function encapsulating the physics of the problem. The integrals are over all the possible solar neutrino energies, $E_{\nu}$, and measurable recoil energies, $E_{R}$. The presence of this double integral makes Eq. 6 a prime candidate for GPU acceleration.
3 Experiments
@@ 3.1 Gaussian Fit with Unknown Mean and Variance
Before applying our sampling methods to the neutrino physics problem, we compare them using a simpler synthetic problem. This is based on a Gaussian fit with unknown mean and standard deviation with a centered parameterization. This is chosen to demonstrate the performance of the various methods using a model that is known to be difficult to sample without reparameterization pmlr-v119-gorinova20a . The likelihood has a similar structure to the well-known Neal’s funnel neal2003slice ; however, this was chosen as a more difficult problem, as Neal’s funnel was efficiently handled by nested sampling as implemented in ultranest 2021JOSS….6.3001B . Our model is given by
@@
We sample from this model using NS and NUTS, both with and without NeuTra. For NS, the number of live points is $1600$ without neural transport and $2400$ with neural transport, as this is found to produce $1.4\times 10^{4}$ effective samples in both cases. We use a sigmoid to transform the parameter space into the unit cube for NS runs. The log-determinant of the transformation Jacobian is summed into the log-density to ensure that the posterior is unaltered by this procedure. For NUTS, each of the $4$ chains we use has a length of 3000 warmup steps and 5000 samples with a target acceptance probability of 0.8. We implement NeuTra with Block Neural Autoregressive Flows 2019arXiv190404676D using two flows with hidden dimensions $[4,4]$. We train the NeuTra model on the ELBO estimated using $30$ points for $5000$ epochs. We run all experiments on an Intel Xeon Gold 6230 CPU. The ground truth evidence and contour integrals are analytically evaluated where possible, and we use torchquad Gomez_torchquad_Numerical_Integration_2021 to compute the remaining integrals via non-stochastic numerical integration; we describe this procedure in greater detail in Appendix C.
Results
Figure 2: Marginalized representations of the inferred posteriors for our synthetic model, Eq. 7, using nested sampling (NS) and the No U-Turn Sampler (NUTS), both with and without neural transport (NeuTra). Left: The $68\%$ and $95\%$ highest posterior density region contours. Middle and Right: Probability density of the samples with $68\%$ equal-tailed intervals indicated as dashed lines. Only $\mu_{1}$ and $\mathcal{C}_{1}$ are shown for clarity; the full corner plot is given in Appendix E. As desired, our results show a high degree of overlap and all closely match the ground truth, derived in Appendix C.
Method Wall Time per ESS [ms] Eval. per ESS Div. per ESS [×$\mathbf{\times 10^{-2}}$]

### v2

# Fast Bayesian Inference for Neutrino Non-Standard Interactions at Dark Matter Direct Detection Experiments

@@ Abstract
Multi-dimensional parameter spaces are commonly encountered in physics theories that go beyond the Standard Model. However, they often possess complicated posterior geometries that are expensive to traverse using techniques traditional to astroparticle physics. Several recent innovations, which are only beginning to make their way into this field, have made navigating such complex posteriors possible. These include GPU acceleration, automatic differentiation, and neural-network-guided reparameterization. We apply these advancements to dark matter direct detection experiments in the context of non-standard neutrino interactions and benchmark their performances against traditional nested sampling techniques when conducting Bayesian inference. Compared to nested sampling alone, we find that these techniques increase performance for both nested sampling and Hamiltonian Monte Carlo, accelerating inference by factors of $\mathord{\sim}100$ and $\mathord{\sim}60$, respectively. As nested sampling also evaluates the Bayesian evidence, these advancements can be exploited to improve model comparison performance while retaining compatibility with existing implementations that are widely used in the natural sciences. Using these techniques, we perform the first scan in the neutrino non-standard interactions parameter space for direct detection experiments whereby all parameters are allowed to vary simultaneously. We expect that these advancements are broadly applicable to other areas of astroparticle physics featuring multi-dimensional parameter spaces.
@@ 1 Introduction
The Standard Model (SM) of particle physics, our best description of the fundamental building blocks of the Universe, is an incomplete description of nature. To better capture reality, theories that go beyond the Standard Model (BSM) are needed [1, 2, 3]. These theories often feature multi-dimensional parameter spaces that must be explored to determine regions that are allowed by the observed data [4, 5, 6, 7, 8, 9]. However, methods for doing so that are traditional to the astroparticle physics community, such as nested sampling (NS), are often inefficient at navigating complicated posterior geometries [10, 11, 12, 13]. This problem is particularly acute in the limit-setting regime, where no significant novel signal is observed and the data is used to restrict the domain of the theory. In this regime, the likelihood is often diffuse and lacks strong modes, rendering it difficult to explore. The problem is exacerbated when the likelihood is expensive to evaluate, as is often true due to the complexity of theoretical frameworks. These issues make it challenging to diagnose convergence failures, incorporate new experimental data, and produce timely studies.
These problems can be tackled using a combination of methods. Leveraging recent powerful numerical computing frameworks, such as JAX [14], and probabilistic programming frameworks, such as numpyro [15, 16], allows us to automatically differentiate likelihood functions and thus sample efficiently using Hamiltonian Monte Carlo (HMC) [17, 18] and the No-U-Turn Sampler (NUTS) [19]. For likelihood functions that are computationally expensive but vectorizable, GPU acceleration can also confer a significant performance boost. Finally, normalizing flows can be used as a reversible transformation to obtain a posterior distribution that is easier to explore in a technique known as neural transport (NeuTra) [20].
The framework of neutrino non-standard interactions (NSI) provides an excellent physics playground on which to benchmark these methods. This framework is a low-energy effective field theory description of BSM neutrino physics, encapsulating the novel interactions that neutrinos could undergo with SM fermions [21, 22, 23, 24, 25, 26, 27]. It features at least $8$ independent parameters: $6$ Wilson coefficients and $2$ angles describing the strength of the new physics interaction with either nucleons or electrons [9, 28]. To explore the NSI landscape fully, we consider dark matter direct detection experiments, which are sensitive to both nuclear recoils (NRs) and electron recoils (ERs) from neutrinos.
While predominantly employed in the search for dark matter, direct detection experiments have improved their sensitivity to the point that they are beginning to approach the ‘neutrino fog’. This marks a region of the WIMP (Weakly Interacting Massive Particle) dark matter parameter space where the rate of NRs due to neutrinos becomes non-negligible, leading to a cumbersome background for dark matter searches [29]. However, this background can be re-imagined as an invaluable signal for direct detection experiments, providing them with the opportunity to probe both SM and BSM neutrino physics. Indeed, the first indications of solar neutrino interactions have recently been observed in direct detection experiments based on dual-phase xenon time projection chambers (TPCs) [30, 31]. ER events are also possible with solar neutrinos and have a different signature from NRs; they can be measured independently, and serve as complementary probes of the neutrino NSI parameter space. These interactions between solar neutrinos and the target material can be used to explore the nature of novel neutrino physics in the NSI framework [9, 32, 33].
@@
2 Sampling Methods
2.1 Nested sampling
Nested sampling (NS) is a family of Monte Carlo methods for calculating integrals over a parameter space [36, 37]. It can also be used for Bayesian computation and has gained popularity for its success in multi-modal problems and its self-tuning feature [38]. NS has many applications in various fields, such as biology [39, 40], linguistics [41, 42] and the physical sciences [43]—including particle physics [44, 32].
@@ 2.1 Nested sampling
$\mathcal{Z}\equiv\int p(\boldsymbol{\theta}|\boldsymbol{X})\,\mathrm{d}\boldsymbol{\theta}\simeq\sum_{i=1}^{N_{\text{iter}}}\Delta V_{i}L_{i}+\frac{V_{\text{end}}}{N}\sum_{j=1}^{N}L_{j},$ (2)
@@
2.2 HMC/NUTS
Hamiltonian Monte Carlo is a powerful sampling algorithm that leverages gradient information to efficiently explore complex probability distributions [17, 18, 45]. HMC offers improved performance compared to traditional Markov chain Monte Carlo (MCMC) methods using a simulation of Hamiltonian dynamics to explore a target parameter space more efficiently than random walks. However, The effectiveness of HMC relies on careful tuning. The NUTS sampler with dual-averaging alleviates this need by adaptively adjusting HMC parameters [46, 47], and it has been widely adopted and implemented across various probabilistic programming frameworks [19, 15, 16, 48, 49].
HMC is extensively used across both the natural and social sciences, and it has been recently employed for analyzing neutrino experiment results [17, 50, 51, 52, 53, 54]. Because HMC is a gradient-based sampling method, it requires the model to either support automatic differentiation or have known derivatives. Interest in it has thus led to new implementations of models using frameworks compatible with automatic differentiation [55, 56].
2.3 Neural Transport
While NS and HMC are better at exploring complex posteriors than simpler methods, such as basic rejection sampling with uniform proposals, sampling efficiency still drops for sufficiently challenging posteriors in multiple dimensions [37, 57]. Under these circumstances, reparameterization can improve the performance of inference algorithms [58], which can be non-trivial for likelihoods that cannot be easily expressed as hierarchical Bayesian models.
One way to remedy this is to use a reversible transformation to map a posterior distribution onto one that is easier to sample, such as a standard normal distribution [20]. Instead of directly sampling from the posterior $p(\boldsymbol{\theta}|\boldsymbol{X})$ using HMC or NS, we can consider a bijective function $\boldsymbol{f}$ that defines the reparameterization of our distribution. This function acts on the reparameterized latent space $\boldsymbol{z}$, such that $\boldsymbol{\theta}=\boldsymbol{f}(\boldsymbol{z})$. We can then instead sample from the distribution
@@ 2.3 Neural Transport
where $|\partial\boldsymbol{f}/\partial\boldsymbol{z}|$ is the determinant of the Jacobian of the parameter transformation. Eq. 3 generally follows from the reparameterization of probability distributions, as shown in A.
Subsequently, the pushforward $\boldsymbol{f}(\boldsymbol{z})$ can be used to obtain the desired samples of $p(\boldsymbol{\theta}|\boldsymbol{X})$. Normalizing flows are a good candidate to parameterize the bijective function, as they can be highly expressive and can be fit by minimizing the Kullback–Leibler (KL) divergence. Using them to reparameterize Bayesian models is known as Neural Transport (NeuTra), as introduced in [20]. There, NeuTra with inverse autoregressive flows was shown to improve sampling performance in problems with difficult posterior distributions. Instead of inverse autoregressive flows, we use block neural autoregressive flows, which demonstrate improved performance over earlier methods [59].
To train a normalizing flow that can transform the posterior distribution of interest into a standard normal distribution, we maximize the evidence lower bound (ELBO) [20, 60], which is equivalent to minimizing the KL divergence [61]. Given $\boldsymbol{f}_{\phi}(\boldsymbol{z})$—the transformation represented by a normalizing flow with parameters $\phi$—the distribution over $\boldsymbol{\theta}$ is
@@
3 Neutrino Non-Standard Interactions and Direct Detection Experiments
3.1 Neutrino Non-Standard Interactions (NSI)
Direct Detection Experiment Solar Neutrino ($\nu$) Target Atom ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\nu}{\color[rgb]{1,1,1}{}_{\alpha}}$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\nu}{\color[rgb]{1,1,1}{}_{\beta}}$ Sun
@@ 3.1 Neutrino Non-Standard Interactions (NSI)
Figure 1: Left: Experimental idea. Solar neutrinos collide with the target atoms in a direct detection experiment, causing either the nucleus or the electrons of the atom to recoil. Right: The neutrino non-standard interactions parameterization, which can be used to model new physics phenomena between neutrinos, neutrons ($n$), protons ($p$), and electrons ($e$) [9]. It is characterized by an overall interaction strength coordinate, $\varepsilon_{\alpha\beta}$, and two angles describing how strong the interaction is with the neutron ($\eta$) and proton vs. electron ($\varphi$).
@@
Neutrino physics that goes beyond the Standard Model can be described using the framework of neutrino non-standard interactions [21, 22, 23, 24, 25, 26, 27]. In this framework, novel neutrino interactions are parametrized within an $8$-dimensional space defined by the coordinates $(\varepsilon_{\alpha\beta},\eta,\varphi)$ [9]. Here, the $6$ parameters $\varepsilon_{\alpha\beta}$ characterize the overall strength and type of interaction. There is one such parameter for each flavor combination $\alpha\beta$ ($ee$, $e\mu$, $e\tau$, etc.), with equal flavor indices meaning that the neutrino type remains the same after the interaction and unequal indices indicating that it changes. The remaining two parameters $(\eta,\varphi)$ are angles describing how strongly the new interaction takes place with protons, neutrons, and electrons. The angle $\eta$ models how strongly the interaction is with the neutron, and the angle $\varphi$ characterizes the strength of the interaction between the electrically charged proton and electron. We visualize the NSI parametrization in Fig. 1 (right)
where the factor of $\sqrt{5}$ is a convention in the field [62, 9]. Except for this factor, this system of equations can be seen to be the projection from the ‘spherical’ NSI parameterization onto a ‘Cartesian’ one.
3.2 Dual-phase Xenon Direct Detection Experiments
Originally leading the hunt for dark matter, dual-phase xenon TPCs are beginning to have sensitivity to solar neutrinos [63]. Two such experiments are XENON1T [64] and PandaX-4T [65]. In these dual-phase xenon TPCs, the energy of the recoiling target due to an interaction with a solar neutrino is converted into scintillation and ionization signals. Measuring both signals allows us to detect low-energy recoils and discriminate between nuclear and electronic events. We provide a schematic of a solar neutrino interaction with a direct detection experiment in Fig. 1 (left). We can count the number of nuclear recoils and/or electron recoilsdetermine whether the signal is consistent with the SM prediction. The NR and ER events constrain the neutrino NSI parameter space in complementary ways, leading to tighter limits or, in the case of a detection, smaller error bars [9, 32].
@@ 3.2 Dual-phase Xenon Direct Detection Experiments
Within the NSI framework, we calculate the number of expected NRs and ERs using the SNuDD Python package [66, 9]. Given a choice of NSI parameters, it computes the number of expected recoils due to solar neutrinos, which is described by the expression [67],
$\frac{\mathrm{d}R}{\mathrm{d}E_{R}}=N_{T}\int_{E_{\nu}^{\mathrm{min}}}\,\frac{\mathrm{d}{\phi_{\nu}}}{\mathrm{d}E_{\nu}}\,\mathrm{Tr}\left[\boldsymbol{\rho}\,\,\frac{\mathrm{d}{\boldsymbol{\zeta}}}{\mathrm{d}E_{R}}\right]\,\mathrm{d}E_{\nu}\,.$ (7)
Here, $N_{T}$ is the number of targets, $\phi_{\nu}$ is the neutrino flux at the source, $\boldsymbol{\rho}$ is the neutrino density matrix at the experiment, $\boldsymbol{\zeta}$ is a generalized scattering cross section, $E_{\nu}$ is the energy of the incident neutrinos, and $E_{\nu}^{\mathrm{min}}$ is the minimum $E_{\nu}$ required to produce a target recoil energy of $E_{R}$.
Both the density matrix and the generalized scattering cross section depend on the full set of NSI parameters. The density matrix dictates the relative neutrino flavor fractions we expect to see on Earth, and it is thus associated with the matter effects stemming from the Mikheyev-Smirnov-Wolfenstein (MSW) effect. Neutrino NSI lead to additional interactions between neutrinos and fermions within the solar medium described by an extra term in the matter Hamiltonian that ultimately alters the density matrix. Similarly, additional potential interactions due to neutrino NSI affect the scattering cross section at the detection point. Full details of both of these effects may be found in Ref. [9].
4 Evaluation of Sampling Methods
@@ 4.1 Gaussian Fit with Unknown Mean and Variance
Before applying our sampling methods to the NSI problem, we compare them using a simpler synthetic problem. This is based on a Gaussian fit with unknown mean and standard deviation with a centered parameterization. This is chosen to demonstrate the performance of the various methods using a model that is known to be difficult to sample without reparameterization [58]. The likelihood has a similar structure to the well-known Neal’s funnel [68]; however, this was chosen as a more difficult problem, as Neal’s funnel was efficiently handled by nested sampling as implemented in ultranest [69]. Our model is given by
@@
We sample from this model using NS and NUTS, both with and without NeuTra. For NS, the number of live points is $1600$ without neural transport and $2400$ with neural transport, as this is found to produce $1.4\times 10^{4}$ effective samples in both cases. We use a sigmoid to transform the parameter space into the unit cube for NS runs. The log-determinant of the transformation Jacobian is summed into the log-density to ensure that the posterior is unaltered by this procedure. For NUTS, each of the $4$ chains we use has a length of 3000 warmup steps and 5000 samples with a target acceptance probability of 0.8. We implement NeuTra with Block Neural Autoregressive Flows [59] using two flows with hidden dimensions $[4,4]$. We train the NeuTra model on the ELBO estimated using $30$ points for $5000$ epochs. We run all experiments on an Intel Xeon Gold 6230 CPU. The ground truth evidence and contour integrals are analytically evaluated where possible, and we use torchquad [70] to compute the remaining integrals via non-stochastic numerical integration; we describe this procedure in greater detail in B.
4.1.1 Results
Figure 2: Marginalized representations of the inferred posteriors for our synthetic model, Eq. 8, using nested sampling (NS) and the No U-Turn Sampler (NUTS), both with and without neural transport (NeuTra). Left: The $68\%$ and $95\%$ highest posterior density region contours. Middle and Right: Probability density of the samples with $68\%$ equal-tailed intervals indicated as dashed lines. Only $\mu_{1}$ and $\mathcal{C}_{1}$ are shown for clarity; the full corner plot is given in D. As desired, our results show a high degree of overlap and all closely match the ground truth, derived in B.
Method Wall Time/ESS [ms] Eval./ESS Div./ESS [×$\mathbf{\times 10^{-2}}$]